Sum of products          sum of squares

X         Y         (X – MX)      (Y – MY)       (X – MX) (Y – MY)                       (X – MX) 2    

4       1                                                                                                

9       3                                                                                               

8       5                                                                          

Sum of products= _____ ?

Sum of Squares= ______ ?

Mx=_____?

My=_____?
Answer #1

x-fly-ỹ (x-xy-)| ON Heren= 3 n Mx = 7 and My = 3 Sum of products (x-Mx)(y- My) = 8 Sum of squares E(x-M_)? =14